OpenStudy (anonymous):

Find the solution to the equation 64^(x – 4) = 4^2x. Use complete sentences to explain the method used to solve this equation.

OpenStudy (anonymous):

4^3^(x-4) = 4^2x 4^3(x-4) = 4^2x 2x = 3(x-4) 2x = 3x - 12 -x = -12 x = 12

OpenStudy (anonymous):

I think you're suppose to use log functions but I'm unsure how to work it out to get the right answer

OpenStudy (anonymous):

Take the log of both sides. (x-4)log(4) 64 = 2x log(4) 4 3(x-4) = 2x Now, solve. x = 12

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

No Problem